OpenStudy (anonymous):

A person is walking in a narrow lane takes 5 steps forward and 3 steps backward, followed again by 5 steps forward and 3 steps backward and so on. each step is 1 m long and requires 1s. How long will it take the person to fall in a pit 13 m away from the start?

OpenStudy (anonymous):

Also find how many steps will it take for him to fall in the pit?

OpenStudy (anonymous):

5m-3m=2m 8sec to go forward 1 time 6times=12m short1m=13m or 48 sec+short1m=49sec

OpenStudy (anonymous):

But the answer in the book is 37seconds

OpenStudy (anonymous):

book is wong

OpenStudy (anonymous):

5 sec f - 3 sec b = 8 sec togo 2m

OpenStudy (anonymous):

2x6=12 so 8x6=48 and we still short 1

OpenStudy (anonymous):

so show the book work or it is miss print it is wrong

OpenStudy (anonymous):

It is correct that the person walks 37m or steps but time taken is 37s as well

OpenStudy (anonymous):

right handx-y=37and left hand x-y=49

OpenStudy (anonymous):

may be I am not thorough with the concepts, but I dont get the solution

OpenStudy (anonymous):

all steps plus or minus are steps stills equals +5- or -3+ equals 8. so if steps equal time go forward and back once still is 8. How did it ever become 6 other than you had to do it 6 times. 6 times 8 plus 1 equals 49. the 2 times 6 plus 1 is 13 but it still only represents how many times you had to go +5 and -3 is still 8 meters. So still my final answer is 49 steps and time. 36 is just the square root of 6, nothing more nothing less. Unless you have some new theory that is my answer. The book is just using the square root trick. Also internet checked, i will take my 12 points. 6 ohter countries came up with the same answer. if you don't believe me take one of your friends and count 5 steps forward 3 steps backwards and see how many steps it takes them to fall in the hole.

OpenStudy (anonymous):

There is another answer on this page: http://cbse-notes.blogspot.in/2012/07/class-11-physics-ch3-motion-in-straight.html

OpenStudy (anonymous):

 30+18+1= 49